How can we experimentally determine how much energy is stored in a corn chip? How does the mass of your chip affect the amount of energy transferred into the water? please provide as much detail as possible. Thank you so much.
Answer #1

To experimentally determine the energy stored in a corn chip follow the following steps-

  1. Do the experimental set up for the energy measurement of Corn chip i.e setting test tube filled with known amount of water over the corn chip. Note down the initial temp of water.
  2. Light the corn chip through a match stick so that it will born.
  3. Now when the chip burn it will definitely release heat to the surrounding.
  4. This heat will the be taken up by the water place above the chip. Thus the temp of the water increase.
  5. Now calculate the increase in temp of water by substracting the final temp after heating and initial temp noted down.
  6. Now the amount of heat absorbed(Q) by the water to increase its temp from T1 to T2 is calculated by

Q = m * C * del T

where Q = amount of heat absorbed by water

m = mass of the water taken

C = specific heat of water = 4.186 J/goC

And del T = change in temp of water = T2 - T1

Now since the heat absorbed by the water is produced from the corn chip burning, this is actually the heat that was stored inside the corn chip.

i.e the above calculated amount of heat absorbed by water = amount of heat stored in chip = Q

2-

Now the amount of heat stored in chip depends upon the mass of the chip through the same formula-

Q = m * C * del T

where Q = amount of heat released from chip

m = mass of the chip

C = specific heat of chip = specific value

And del T = change in temp of chip.

This formula shows that the amount of heat released is directly proportional to the mass of the chip. Thus as the mass of the chip increases, the amount of heat released will also increase. Thus the amount of heat absorbed by the water also increases.